# CountingFruit: Language-Guided 3D Fruit Counting with Semantic Gaussian Splatting

### v2

@@ Abstract
Accurate 3D fruit counting in orchards is challenging due to heavy occlusion, semantic ambiguity between fruits and surrounding structures, and the high computational cost of volumetric reconstruction. Existing pipelines often rely on multi-view 2D segmentation and dense volumetric sampling, which lead to accumulated fusion errors and slow inference. We introduce FruitLangGS, a language-guided 3D fruit counting framework that reconstructs orchard-scale scenes using an adaptive-density Gaussian Splatting pipeline with radius-aware pruning and tile-based rasterization, enabling scalable 3D representation. During inference, compressed CLIP-aligned semantic vectors embedded in each Gaussian are filtered via a dual-threshold cosine similarity mechanism, retrieving Gaussians relevant to target prompts while suppressing common distractors (e.g., foliage), without requiring retraining or image-space masks. The selected Gaussians are then sampled into dense point clouds and clustered geometrically to estimate fruit instances, remaining robust under severe occlusion and viewpoint variation. Experiments on nine different orchard-scale datasets demonstrate that FruitLangGS consistently outperforms existing pipelines in instance counting recall, avoiding multi-view segmentation fusion errors and achieving up to 99.2% recall on Fuji-SfM orchard dataset. Ablation studies further confirm that language-conditioned semantic embedding and dual-threshold prompt filtering are essential for suppressing distractors and improving counting accuracy under heavy occlusion. Beyond fruit counting, the same framework enables prompt-driven 3D semantic retrieval without retraining, highlighting the potential of language-guided 3D perception for scalable agricultural scene understanding.
@@ 3.2 Language-Conditioned Semantic Embedding
To reduce memory consumption and improve runtime efficiency, we compress these 512512512-dimensional CLIP vectors into a lower-dimensional latent space. We train a lightweight scene-specific autoencoder $\Psi\circ E$ using the loss:
@@
Dataset Metric NeRF FruitNeRF Taming3DGS 3DGS FruitLangGS (Ours)
@@
FPS $\uparrow$ $<$0.5 1.44 164.89 128.36 807
3.3 Prompt-Based Filtering and Counting
=$\mathcal{M}=\left\{i\mid s_{i}^{+}>\tau_{\text{pos}}\;\land\;s_{i}^{-}<\tau_{\text{neg}}\right\},$ (11)
$N_{i}\propto\det(\Sigma_{i})^{\frac{1}{2}}\cdot\alpha_{i},$ (12)
$\hat{N}_{\text{fruit}}=\sum_{j=1}^{r}\gamma(C_{j}),$ (13)
4 Experiments and Results
@@
Pfuji-Size_Orch2018 1075 (73.9%) 1238 (85.1%) 1254 (86.2%) 1336 (91.8%) 1455
Pfuji-Size_Orch2020 1160 (79.7%) 1222 (84.0%) 1201 (82.5%) 1329 (91.3%) 1455
@@
4.1 Datasets and Implementation Details
  • FruitNeRF Benchmark (3Tree and Fuji-SfM) [9]: Includes (1) 3Tree—three high-res Fuji trees ($\sim$350 posed RGBs each, 4000×60004000$\times$ 60004000 × 6000) with instance-level ground-truth; and (2) Fuji-SfM—eleven trees reconstructed via SfM, with only tree-level counts for large-scale aggregation.
All experiments are conducted on a workstation with dual NVIDIA RTX 4090 GPUs and 128 GB RAM. The pipeline is implemented in PyTorch with full GPU acceleration and operates offline. For each sequence, 3D Gaussian parameters are trained for 30,000 iterations with a learning rate of $1.6\times 10^{-3}$ and exponential decay. Images are resized to 1000×15001000$\times$ 15001000 × 1500, and rendering uses spherical harmonics of order $l=3$. Progressive densification, opacity pruning, and geometric regularization are applied. Semantic filtering, sampling, and clustering hyperparameters are listed in Table 1.
@@ 4.1 Datasets and Implementation Details
Pfuji-Size_Orch2018 1336 / 1455 (91.8%) 1188 / 1455 (81.6%) 1254 / 1455 (86.2%)
Pfuji-Size_Orch2020 1329 / 1455 (91.3%) 1174 / 1455 (80.7%) 1201 / 1455 (82.5%)
@@ 4.3 Instance Counting Accuracy
On the Tree03, our method identifies 287 out of 291 fruits, achieving 98.6% recall, 31.9% higher than 3DGS+Ours and 7.9% higher than FruitNeRF. On Tree01 and Tree02, our method performs comparably to FruitNeRF, achieving 96.6% and 97.3% recall, respectively. Across the large-scale Fuji-SfM set, FruitLangGS detects 1443 of 1455 fruits (99.2% recall), matching FruitNeRF (1459 detections) but without inflated oversegmentation. Furthermore, on Pfuji-Size_Orch2018 and Orch2020 subsets, our method reaches recall rates of 91.8% and 91.3%, outperforming all baselines. For the Galafab-west, Schnico-Red-east, and Schniga-Schnico-west datasets, where per-tree annotations are unavailable, we report estimated counts from full-scene reconstructions. These results affirm the generalizability of FruitLangGS to unseen domains with complex occlusions and variable capture conditions.

### v3

@@ Abstract
Accurate 3D fruit counting in orchards is challenging due to heavy occlusion, semantic ambiguity between fruits and surrounding structures, and the high computational cost of volumetric reconstruction. Existing pipelines often rely on multi-view 2D segmentation and dense volumetric sampling, which lead to accumulated fusion errors and slow inference. We introduce FruitLangGS, a language-guided 3D fruit counting framework that reconstructs orchard-scale scenes using an adaptive-density Gaussian Splatting pipeline with radius-aware pruning and tile-based rasterization, enabling scalable 3D representation. During inference, compressed CLIP-aligned semantic vectors embedded in each Gaussian are filtered via a dual-threshold cosine similarity mechanism, retrieving Gaussians relevant to target prompts while suppressing common distractors (e.g., foliage), without requiring retraining or image-space masks. The selected Gaussians are then sampled into dense point clouds and clustered geometrically to estimate fruit instances, remaining robust under severe occlusion and viewpoint variation. Experiments on nine different orchard-scale datasets demonstrate that FruitLangGS consistently outperforms existing pipelines in instance counting recall, avoiding multi-view segmentation fusion errors and achieving up to 99.7% recall on Pfuji-Size_Orch2018 orchard dataset. Ablation studies further confirm that language-conditioned semantic embedding and dual-threshold prompt filtering are essential for suppressing distractors and improving counting accuracy under heavy occlusion. Beyond fruit counting, the same framework enables prompt-driven 3D semantic retrieval without retraining, highlighting the potential of language-guided 3D perception for scalable agricultural scene understanding.
@@ 3.2 Language-Conditioned Semantic Embedding
To reduce memory consumption and improve runtime efficiency, we compress these $512$-dimensional CLIP vectors into a lower-dimensional latent space. We train a lightweight scene-specific autoencoder $\Psi\circ E$ using the loss:
@@
3.3 Prompt-Based Filtering and Counting
$\mathcal{M}=\left\{i\mid s_{i}^{+}>\tau_{\text{pos}}\;\land\;s_{i}^{-}<\tau_{\text{neg}}\right\},$ (11)
$N_{i}\propto\det(\Sigma_{i})^{\frac{1}{2}}\cdot\alpha_{i},$ (12)
$\hat{N}_{\text{fruit}}=\sum_{j=1}^{r}\gamma(C_{j}),$ (13)
4 Experiments and Results
@@
Dataset Metric NeRF FruitNeRF Taming3DGS 3DGS FruitLangGS (Ours
@@
FPS $\uparrow$ <<0.5 1.44 164.89 128.36 807
@@
Pfuji-Size_Orch2018 224 (73.9%) 294 (97.0%) 261 (86.1%) 302 (99.7%) 303
Pfuji-Size_Orch2020 249 (79.8%) 293 (93.9%) 257 (82.4%) 307 (98.4%) 312
@@
Pfuji-Size_Orch2018 302 / 303 (99.7%) 247 / 303 (81.5%) 261 / 303 (86.1%)
Pfuji-Size_Orch2020 307 / 312 (98.4%) 252 / 312 (80.8%) 257 / 312 (82.4%)
4.1 Datasets and Implementation Details
  • FruitNeRF Benchmark (3Tree and Fuji-SfM) [9]: Includes (1) 3Tree—three high-res Fuji trees ($\sim$350 posed RGBs each, $4000\times 6000$) with instance-level ground-truth; and (2) Fuji-SfM—eleven trees reconstructed via SfM, with only tree-level counts for large-scale aggregation.
All experiments are conducted on a workstation with dual NVIDIA RTX 4090 GPUs and 128 GB RAM. The pipeline is implemented in PyTorch with full GPU acceleration and operates offline. For each sequence, 3D Gaussian parameters are trained for 30,000 iterations with a learning rate of $1.6\times 10^{-3}$ and exponential decay. Images are resized to $1000\times 1500$, and rendering uses spherical harmonics of order $l=3$. Progressive densification, opacity pruning, and geometric regularization are applied. Semantic filtering, sampling, and clustering hyperparameters are listed in Table 1.
@@ 4.3 Instance Counting Accuracy
On the Tree03, our method identifies 287 out of 291 fruits, achieving 98.6% recall, 31.9% higher than 3DGS+Ours and 7.9% higher than FruitNeRF. On Tree01 and Tree02, our method performs comparably to FruitNeRF, achieving 96.6% and 97.3% recall, respectively. Across the large-scale Fuji-SfM set, FruitLangGS detects 1443 of 1455 fruits (99.2% recall), matching FruitNeRF (1459 detections) but without inflated oversegmentation. Furthermore, on Pfuji-Size_Orch2018 and Orch2020 subsets, our method reaches recall rates of 99.7% and 98.4%, outperforming all baselines. For the Galafab-west, Schnico-Red-east, and Schniga-Schnico-west datasets, where per-tree annotations are unavailable, we report estimated counts from full-scene reconstructions. These results affirm the generalizability of FruitLangGS to unseen domains with complex occlusions and variable capture conditions.